[Brahms-l] Shift report 20050316 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 16 2005 - 08:00:49 EST
Shifter leader was SB
Shifters were 


1. Two stores. First one from 00:45-05:30, Second one from 06:30-
The rate for trig2 ~ 4, going slowly
----------------------------------------------
2. Run No.   FS@8 deg @A1219  MRS@45 deg @A700
----------------------------------------------
   13808     trig2(0.5k)          trig3(14k)
   13809     trig2(34k)           trig3(743k)
----------------------------------------------
3. No detector problems observed

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 16 08:08:52 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 16 2005 - 08:09:01 EST